Refer to the information provided in Toble 8.5 below to answer the questions that follow Table 8.5 S million 2,800 3,600 4,400 Refer to Table 8.5. The equilibrium level of aggregate output equals O $3,000 million O S4.000 milion. $6.000 million. ! Question 22 1 pts lf C . 100 + 0.8Y and I-50, then the equilibrium level of income is 1875. O 375 O 600 O 750.
Answer #1

1. At equilibrium level of income : Aggregate output (Y) = Aggregate expenditure (AE)

AE = C+I

Aggregate output (Y) C I AE
3000 2000 1600 3600
4000 2800 1600 4400
5000 3600 1600 5200
6000 4400 1600 6000
7000 5200 1600 6800

The equilibrium level of aggregate output equals $6000 million. Hence,option(D) is correct.

2. C= 100+ 0.8Y

I =50

AE = C+I

= 100+ 0.8Y + 50

AE = 150 + 0.8Y

Ar equilibrium : Y =AE

Y = 150+ 0.8Y

Y- 0.8Y = 150

0.2Y= 150

Y =750

The equilibrium level of income = 750.

Hence,option(D) is correct.